# AC-PKAN: Attention-Enhanced and Chebyshev Polynomial-Based Physics-Informed Kolmogorov–Arnold Networks

### v1

@@ 3.4 Residual-and-Gradient Based Attention
In the canonical PINN formulation, the loss is split into an unlabeled PDE-residual term $\mathcal{L}_{r}$ and a labeled term $\mathcal{L}_{d}$ that enforces boundary/initial constraints and matches available data samples. To improve optimization efficiency and accuracy—while counteracting the loss imbalance introduced by Chebyshev bases—we introduce Residual–Gradient Attention (RGA), an adaptive scheme that rescales each loss term according to its residual magnitude and corresponding gradient norm. This approach ensures balanced and efficient optimization, particularly addressing challenges with boundary and initial condition losses.
@@
Residual-Based Attention (RBA) dynamically amplifies loss terms with the largest point-wise residuals, assigning a tensor of weights $w_{i,j}^{\mathrm{RBA}}$ to each loss component $\mathcal{L}_{i}$ ($i\in\{r,d\}$) at location $j$ (anagnostopoulos2024residual, ):
@@
where $\eta$ is the RBA learning rate and $\max_{j}\lvert\mathcal{L}_{i,j}\rvert$ normalizes by the maximal residual. As a lightweight, point-wise weighting scheme, RBA complements the Cheby1KAN layer—which excels at capturing strong nonlinearity and complex distributions but can suffer from slow or unstable convergence—by embedding a self-adjusting feedback loop. This synergy alleviates numerical optimization difficulties and enhances global convergence efficiency.
@@
GRA dynamically adjusts weights based on gradient norms of different loss components, promoting balanced training. As a scalar applied to one entire loss term, GRA addresses the imbalance where gradient norms of the PDE residual loss significantly exceed those of the data fitting loss (wang2021understanding, ), which can lead to pathological gradient flow issues (wang2022and, ; fang2023ensemble, ). Our mechanism smooths weight adjustments, preventing the network from overemphasizing residual loss terms and neglecting other essential physical constraints, thus enhancing convergence and stability.
@@
To equilibrate the magnitudes of GRA and RBA weights, we apply a logarithmic transformation to the GRA weights when incorporating them into the loss terms, while retaining their original form during weight updates. This preserves the direct relationship between weights and gradient information, ensuring sensitivity to discrepancies between residual and data gradients. The logarithmic transformation mitigates magnitude disparities, preventing imbalances among loss terms. It enables GRA weights to adjust more rapidly when discrepancies are minor and ensures stable updates when discrepancies are substantial. The coefficient $\lambda^{\text{GRA}}$ not only attains excessively large values in scale but also exhibits a broad range of variation. In the training process, $\lambda^{\text{GRA}}$ rapidly increases from zero to very large values, demonstrating a wide dynamic range which is shown in Figure 4 in Appendix I. The logarithmic transformation significantly constrains this range; without it, the model cannot accommodate drastic changes in $\lambda^{\text{GRA}}$, and rigid manual scaling factors further exacerbate the imbalance among loss terms, ultimately causing training failure. For details on the effect of logarithmic transformation in the RGA module, see Appendix D.
@@
This formulation reweights the residual loss based on its magnitude and adjusts the boundary and initial condition losses according to both their magnitudes and gradient norms, promoting balanced and focused training through a dual attention mechanism. The whole algorithmic details are provided in algorithm 1 in Appendix A.
@@ 4 Experiments
Goal. Our empirical study highlights three principal strengths of AC-PKAN: (1) its internal architecture delivers powerful symbolic representation and function-approximation capabilities even in the absence of the RGA loss-weighting scheme; (2) it significantly improves generalization abilities and mitigates failure modes compared to PINNs and other KAN variants; and (3) it achieves superior performance in complex real-world engineering environments. To substantiate our claims, we assemble three task suites (nine benchmarks in total) and benchmark 12 representative architectures—including PINN, PINNsFormer, KAN, fKAN and other variants. For detailed methodology underpinning our first objective, see the Supplementary Experiment in Appendix G.1. While operator-learning frameworks that depend on extensive volumes of labeled data have lately dominated SciML (lu2019deeponet, ; li2020fourier, ; li2020neural, ; tripura2022wavelet, ; calvello2024continuum, ), we follow the customary scope of PINN-refinement studies—which target unsupervised or weakly-supervised regimes—and  (jagtap2020extended, ; yu2022gradient, ; zhang2025bo, ; hou2024reconstruction, ; li2023adversarial, )therefore do not report operator-learning baselines, whose reliance on dense labelled data renders them orthogonal to our setting. The experimental setup was inspired by methodologies in (chebykan2024, ; hao2023pinnacle, ; wang2024kolmogorov, ; zhao2023pinnsformer, ; wang2023expert, ). In all experiments, the best results are highlighted in bold italics, and the second-best results in bold.
@@ 4.1 Mitigating Failure Modes in PINNs
We assessed the AC-PKAN model on five complex PDEs known as PINN failure modes—the 1D-Wave PDE, 1D-Reaction PDE, 2D Navier–Stokes Flow around a Cylinder, 1D Convection-Diffusion-Reaction and 2D Navier–Stokes Lid-driven Cavity Flow (mojgani2022lagrangian, ; daw2022mitigating, ; krishnapriyan2021characterizing, )—to demonstrate its superior generalization ability compared to other PINN variants. In these cases, optimization often becomes trapped in local minima, leading to overly smooth approximations that deviate from true solutions.
Evaluation results are summarized in Table 1, with detailed PDE formulations and setups in Appendix H. Prediction for AC-PKAN are shown in Figure  2 and additional plots including the analysis of loss landscapes are in Appendix I.
@@
4.2 PDEs in Complex Engineering Environments
We further evaluated AC-PKAN across three challenging scenarios: heterogeneous environments, complex geometric boundary conditions, and three-dimensional spatial point clouds. Literature indicates that PINNs encounter difficulties with heterogeneous problems due to sensitivity to material properties (aliakbari2023ensemble, ), significant errors near boundary layers (piao2024domain, ), and convergence issues (sumanta2024physics, ). Additionally, original KANs perform poorly with complex geometries (wang2024kolmogorov, ). The sparsity, irregularity, and high dimensionality of unstructured 3D point cloud data hinder PINNs from effectively capturing spatial features, resulting in suboptimal training performance (chen2022deepurbandownscale, ). We applied AC-PKAN to solve Poisson equations within these contexts.
Detailed PDE formulations are in Appendix H, and detailed experimental results are illustrated in Appendix I. Summarized in Table 2 and partially shown in Figure 2, the results indicate that AC-PKAN consistently achieves the best or second-best performance. It demonstrates superior potential in solving heterogeneous problems without subdomain division and exhibits promising application potential in complex geometric boundary problems where most models fail.
4.3 Ablation Study
Module importance.
Model rMAE rRMSE
AC-PKAN 0.0011 0.0011
AC-PKAN (no GRA) 0.0779 0.0787
AC-PKAN (no RBA) 0.0494 0.0500
AC-PKAN (no RGA) 0.4549 0.4488
AC-PKAN (no Wavelet) 0.0045 0.0046
AC-PKAN (no Encoder) 0.0599 0.0584
AC-PKAN (no MLPs) 1.0422 1.0246
Transferability of RGA.
@@ 5 Conclusion
We introduced AC-PKAN, a novel framework that enhances PINNs by integrating Cheby1KAN with traditional MLPs and augmenting them with internal and external attention mechanisms. This improves the model’s ability to capture complex patterns and dependencies, resulting in superior performance on challenging PDE tasks.
Appendix A Pseudocode of the Algorithm
Algorithm 1 Implementation of the RGA Mechanism
Data: Model parameters $\theta$, total number of parameters $P$, learning rate $\alpha$, hyperparameters $\eta$, $\beta_{w}$, $\epsilon$
Initialization: $w_{r,d}^{\text{RBA}}\leftarrow 0$, $\lambda_{d}^{\text{GRA}}\leftarrow 1$
1:for each training iteration do
2:    Compute gradients:
3:    Update RBA weights for each data point $j$:
$w_{i,j}^{\text{RBA}}\leftarrow(1-\eta)w_{i,j}^{\text{RBA}}+\eta\left(\dfrac{|\mathcal{L}_{i,j}|}{\max_{j}|\mathcal{L}_{i,j}|}\right),\quad i\in\{r,d\}$
4:    Compute gradient norms:
$G_{r}^{\max}\leftarrow\max_{p}\left\|\nabla_{\theta_{p}}\mathcal{L}_{r}\right\|,\quad\overline{G}_{i}\leftarrow\dfrac{1}{P}\sum_{p=1}^{P}\left\|\nabla_{\theta_{p}}\mathcal{L}_{i}\right\|,\quad i\in\{d\}$
5:    Update GRA weights:
$\hat{\lambda}_{i}\leftarrow\dfrac{G_{r}^{\max}}{\epsilon+\overline{G}_{i}},\quad\lambda_{i}^{\text{GRA}}\leftarrow(1-\beta_{w})\lambda_{i}^{\text{GRA}}+\beta_{w}\hat{\lambda}_{i},\quad\lambda_{i}^{\text{GRA}}\leftarrow\max\left(e+\epsilon,\lambda_{i}^{\text{GRA}}\right),\quad i\in\{d\}$
6:    Compute total loss:
$\mathcal{L}_{\text{RGA}}\leftarrow\lambda_{r}w_{r}^{\text{RBA}}\mathcal{L}_{r}+\sum_{i\in\{d\}}\lambda_{i}w_{i}^{\text{RBA}}\log\left(\lambda_{i}^{\text{GRA}}\right)\mathcal{L}_{i}$
7:    Update model parameters:
$\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{RGA}}$
8:end for
Algorithm 2 Internal AC-PKAN Forward Pass
Data: Input data $\mathbf{x}$, Cheby1KAN layer parameters, Wavelet activation function parameters
1:Input embedding:
2:Compute representations:
$\mathbf{U}\leftarrow\text{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{U}+\mathbf{b}_{U}),\quad\mathbf{V}\leftarrow\text{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{V}+\mathbf{b}_{V})$
3:Initialize attention:
$\alpha^{(0)}\leftarrow\mathbf{U}$
4:for $l=1$ to $L$ do
5:    Update attention:
6:
$\mathbf{H}^{(l)}\leftarrow\text{LayerNorm}\left(\text{Cheby1KANLayer}\left(\alpha^{(l-1)}\right)\right)$
$\alpha^{(l)}_{0}\leftarrow\mathbf{H}^{(l)}+\alpha^{(l-1)}$
$\alpha^{(l)}\leftarrow(1-\alpha^{(l)}_{0})\odot\mathbf{U}+\alpha^{(l)}_{0}\odot\left(\mathbf{V}+1\right)$
7:end for
8:Output prediction:
Appendix B Mathematical Proofs
B.1 Proof of Theorem 3.1
Lemma B.1.
@@ Appendix B Mathematical Proofs
By the properties of subspace dimensions, it follows from (23) that
@@
Therefore, from (25) and (26), it follows that
@@
Combining (24) and (27), we obtain
@@
Applying the previously established result (28), we obtain
@@
By the inductive hypothesis (30), we have
@@
Therefore, substituting (33) into (32), we obtain
@@
By induction, the inequality (29) holds for all $n\geq 2$. This completes the proof. ∎
Theorem B.2 (Single Cheb1KAN Layer Rank Constraint).
@@
To conclude that $N+1$ Chebyshev polynomials $T_{0}(x),T_{1}(x),\cdots,T_{N}(x)$ are linearly independent on interval $[-1,1]$, we assume the opposite: there exist constants $c_{0},c_{1},\cdots,c_{N}$ such that:
$\forall x\in[-1,1],\quad c_{0}T_{0}(x)+c_{1}T_{1}(x)+\cdots+c_{N}T_{N}(x)=0$ (38)
Since each $T_{k}(x)$ are a set of k-degree orthogonal polynomials according to  8, the left side is a polynomial of degree at most $N$. A non-zero polynomial of degree $N$ can have at most $N$ roots. However, the equation holds for all $x$ in $[-1,1]$, which is an infinite set of points. Therefore, the polynomial must be the zero polynomial, implying $c_{0}=c_{1}=\cdots=c_{N}=0$.
@@
The key to this upper bound is that the basis function expansions for different input vectors are independent. Based on Equation  36, although the output dimension $d_{l+1}$ may be much smaller than $d_{l}\cdot(N+1)$, the final column space dimension is constrained by the following two factors:
@@
B.2 Proof of Theorem 3.2
Theorem B.3 (Nonlinear Normalization Effect).

### v2

@@ 3.4 Residual-and-Gradient Based Attention
In the canonical PINN formulation, the loss is split into an unlabeled PDE-residual term $\mathcal{L}_{r}$ and a labeled term $\mathcal{L}_{d}$ that enforces boundary/initial constraints and matches available data samples. To improve optimization efficiency and accuracy and to correct the loss imbalance introduced by Chebyshev bases, we propose Residual Gradient Attention (RGA), an adaptive mechanism that rescales each loss term according to its residual magnitude and its gradient norm. This approach ensures balanced and efficient optimization, particularly addressing challenges with boundary and initial condition losses.
@@
Residual‐Based Attention (RBA) dynamically amplifies loss terms with the largest point‐wise residuals, assigning a tensor of weights $w_{i,j}^{\mathrm{RBA}}$ to each loss component $\mathcal{L}_{i}$ ($i\in\{r,d\}$) at location $j$ (Anagnostopoulos et al., 2024):
@@
where $\eta$ is the RBA learning rate and $\max_{j}\lvert\mathcal{L}_{i,j}\rvert$ normalizes by the maximal residual. Residual-Based Attention (RBA) is a lightweight, pointwise weighting scheme that complements the Cheby1KAN layer. Cheby1KAN captures strong nonlinear structure and complex distributions but can exhibit slow or unstable convergence. RBA inserts a self-adjusting feedback loop that reweights local training signals according to residual statistics, focusing learning on poorly fitted locations and reducing the influence of noisy or saturated terms. This synergy alleviates numerical optimization difficulties and enhances global convergence efficiency.
@@
GRA dynamically adjusts weights based on gradient norms of different loss components, promoting balanced training. As a scalar applied to one entire loss term, GRA addresses the imbalance where gradient norms of the PDE residual loss significantly exceed those of the data fitting loss (Wang et al., 2021), which can lead to pathological gradient flow issues (Wang et al., 2022b; Fang et al., 2023). Our mechanism smooths weight adjustments, preventing the network from overemphasizing residual loss terms and neglecting other essential physical constraints, thus enhancing convergence and stability.
@@
Algorithm 2 Implementation of the RGA Mechanism
Data: Model parameters $\theta$, total number of parameters $P$, learning rate $\alpha$, hyperparameters $\eta,\ \beta_{w},\ \epsilon$.
Initialization: $w_{r,d}^{\mathrm{RBA}}\leftarrow 0,\quad\lambda_{d}^{\mathrm{GRA}}\leftarrow 1$.
1:for each training iteration do
2:  Compute gradients:
3:  Update RBA weights for each data point $j$:
$w_{i,j}^{\mathrm{RBA}}\leftarrow(1-\eta)w_{i,j}^{\mathrm{RBA}}+\eta\left(\dfrac{|\mathcal{L}_{i,j}|}{\max_{j}|\mathcal{L}_{i,j}|}\right),\quad i\in\{r,d\}$
4:  Compute gradient norms:
$G_{r}^{\max}\leftarrow\max_{p}\|\nabla_{\theta_{p}}\mathcal{L}_{r}\|,\quad\overline{G}_{i}\leftarrow\dfrac{1}{P}\sum_{p=1}^{P}\|\nabla_{\theta_{p}}\mathcal{L}_{i}\|,\quad i\in\{d\}$
5:  Update GRA weights:
$\hat{\lambda}_{i}\leftarrow\dfrac{G_{r}^{\max}}{\epsilon+\overline{G}_{i}},\quad\lambda_{i}^{\mathrm{GRA}}\leftarrow(1-\beta_{w})\lambda_{i}^{\mathrm{GRA}}+\beta_{w}\hat{\lambda}_{i},\quad\lambda_{i}^{\mathrm{GRA}}\leftarrow\max(\epsilon,\,\lambda_{i}^{\mathrm{GRA}}),\quad i\in\{d\}$
6:  Compute total loss:
$\mathcal{L}_{\mathrm{RGA}}\leftarrow\lambda_{r}\,w_{r}^{\mathrm{RBA}}\,\mathcal{L}_{r}\;+\;\sum_{i\in\{d\}}\lambda_{i}\,w_{i}^{\mathrm{RBA}}\,\log(\lambda_{i}^{\mathrm{GRA}})\,\mathcal{L}_{i}$
7:  Update model parameters:
$\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathrm{RGA}}$
8:end for
@@
To equilibrate the magnitudes of GRA and RBA weights, we apply a logarithmic transformation to the GRA weights when incorporating them into the loss terms, while retaining their original form during weight updates. This preserves the direct relationship between weights and gradient information, ensuring sensitivity to discrepancies between residual and data gradients. The logarithmic transformation mitigates magnitude disparities, preventing imbalances among loss terms. It enables GRA weights to adjust more rapidly when discrepancies are minor and ensures stable updates when discrepancies are substantial. The coefficient $\lambda^{\text{GRA}}$ not only attains excessively large values in scale but also exhibits a broad range of variation. In the training process, $\lambda^{\text{GRA}}$ rapidly increases from zero to very large values, demonstrating a wide dynamic range which is shown in Figure 4 in Appendix F. The logarithmic transformation significantly constrains this range; without it, the model cannot accommodate drastic changes in $\lambda^{\text{GRA}}$, and rigid manual scaling factors further exacerbate the imbalance among loss terms, ultimately causing training failure.
@@
This formulation reweights the residual loss based on its magnitude and adjusts the boundary and initial condition losses according to both their magnitudes and gradient norms, promoting balanced and focused training through a dual attention mechanism.
@@ 4 Experiments
Goal. Our empirical study highlights three principal strengths of AC-PKAN: (1) its internal architecture delivers powerful symbolic representation and function-approximation capabilities; (2) it significantly improves generalization abilities and mitigates failure modes compared to PINNs and other KAN variants; and (3) it achieves superior performance in complex real-world engineering environments. We evaluate our method on three task suites comprising nine benchmarks and compare it to 12 representative architectures, including PINN, PINNsFormer, KAN, and fKAN. Although operator learning frameworks that rely on large volumes of labeled data have recently dominated SciML (Lu et al., 2019; Li et al., 2020a; b; Tripura and Chakraborty, 2022; Calvello et al., 2024), our work remains within the established remit of PINN refinement studies, which focus on unsupervised or weakly supervised settings (Jagtap and Karniadakis, 2020; Yu et al., 2022; Zhang et al., 2025a; Hou et al., 2024; Li et al., 2023). Operator-learning methods typically require dense supervision and therefore address a different set of trade-offs compared to physics-informed approaches. Nevertheless, we include a head-to-head comparison of operator-learning and PINN-variant baselines under sparse-data conditions in Section˜C.3. The experimental setup was inspired by methodologies in (SS, 2024; Hao et al., 2023; Wang et al., 2024b; Zhao et al., 2023; Wang et al., 2023). In all experiments, the best results are highlighted in bold italics, and the second-best results in bold. For the formal definitions of the evaluation metrics and detailed descriptions of the experimental setup, please refer to appendix˜E.
4.1 Complex Function Fitting
We evaluated the AC-PKAN Simplified model, which retains only the internal architecture, on a challenging function interpolation benchmark and compared its performance to a PINN implemented as an MLP, the original KAN, and several KAN variants. Detailed experimental setups and results are provided in Appendices E and F.
As shown in Figure 3(b), the AC-PKAN Simplified model converges more rapidly than MLPs, KAN, and most KAN variants, achieving lower final losses. While Cheby2KAN and FourierKAN demonstrate faster convergence, our model produces smoother fitted curves and exhibits greater robustness to noise, effectively preventing overfitting in regions with high-frequency variations.Performance metrics are presented in Table 3(a).
Model rMAE rMSE Loss
Cheby1KAN 0.0179 0.0329 0.0068
Cheby2KAN 0.0189 0.0313 0.0079
MLP 0.0627 0.1250 0.1410
AC-PKAN_s 0.0177 0.0311 0.0081
KAN 0.0145 0.0278 0.0114
rKAN 0.0458 0.0783 0.1867
fKAN 0.0858 0.1427 0.1722
FastKAN 0.0730 0.1341 0.1399
FourierKAN 0.0211 0.0353 0.0063
(a) Comparison of test rMAE, rMSE, and training Loss
(b) Convergence Comparison of Nine Models
@@ 4.2 Mitigating Failure Modes in PINNs
We assessed the AC-PKAN model on five complex PDEs known as PINN failure modes—the 1D-Wave PDE, 1D-Reaction PDE, 2D Navier–Stokes Flow around a Cylinder, 1D Convection-Diffusion-Reaction and 2D Navier–Stokes Lid-driven Cavity Flow (Mojgani et al., 2022; Daw et al., 2022; Krishnapriyan et al., 2021)—to demonstrate its superior generalization ability compared to other PINN variants. In these cases, optimization often becomes trapped in local minima, leading to overly smooth approximations that deviate from true solutions.
Evaluation results are summarized in Table 1, with detailed PDE formulations and setups in Appendix E. Prediction for AC-PKAN are shown in Figure 2 and additional plots including the analysis of loss landscapes are in Appendix F.
@@
4.3 PDEs in Complex Engineering Environments
We further evaluated AC-PKAN across three challenging scenarios: heterogeneous environments, complex geometric boundary conditions, and three-dimensional spatial point clouds. Literature indicates that PINNs encounter difficulties with heterogeneous problems due to sensitivity to material properties (Aliakbari et al., 2023), significant errors near boundary layers (Piao et al., 2024), and convergence issues (Sumanta et al., 2024). Additionally, original KANs perform poorly with complex geometries (Wang et al., 2024b). The sparsity, irregularity, and high dimensionality of unstructured 3D point cloud data hinder PINNs from effectively capturing spatial features, resulting in suboptimal training performance (Chen et al., 2022). We applied AC-PKAN to solve Poisson equations within these contexts.
Model rMAE rRMSE
AC-PKAN 0.0011 0.0011
AC-PKAN (no GRA) 0.0779 0.0787
AC-PKAN (no RBA) 0.0494 0.0500
AC-PKAN (no RGA) 0.4549 0.4488
AC-PKAN (no Wavelet) 0.0045 0.0046
AC-PKAN (no Encoder) 0.0599 0.0584
AC-PKAN (no MLPs) 1.0422 1.0246
Detailed PDE formulations are in Appendix E, and detailed experimental results are illustrated in Appendix F. Summarized in Table 2 and partially shown in Figure 2, the results indicate that AC-PKAN consistently achieves the best or second-best performance. It demonstrates superior potential in solving heterogeneous problems without subdomain division and exhibits promising application potential in complex geometric boundary problems where most models fail.
4.4 Ablation Study
Module importance.
Transferability of RGA.
Effect of Logarithmic Transformation in the RGA Module.
In this ablation study, we investigated the impact of removing the logarithmic transformation in the RGA module across five PDE experimental tasks. To compensate for the absence of the logarithmic scaling, we adjusted the scaling factors to smaller values. Specifically, we employed the original RGA design to pre-train the models for several epochs, during which very large values of $\lambda^{\text{GRA}}$ were obtained. We fix the PDE residual scale to one and set the data loss scales, including boundary and initial condition terms, to be inversely proportional to the current scale of $\lambda^{\mathrm{GRA}}$, which keeps the different loss contributions at comparable magnitudes.
Equation Without Log With Log
rMAE rRMSE rMAE rRMSE
2D NS Cylinder 532.2411 441.0240 0.2230 0.2182
1D Wave 0.7686 0.7479 0.0011 0.0011
1D Reaction 2.2348 2.2410 0.0375 0.0969
Heterogeneous Problem 10.0849 9.6492 0.1063 0.1817
Complex Geometry 164.4283 158.7840 0.5452 0.5896
Integration with Other External Learning Strategies for Enhanced Performance of AC-PKAN.
Model rMAE rRMSE
AC-PKAN + NTK 0.0009 0.0009
PINNs + NTK 0.1397 0.1489
PINNsFormer + NTK 0.0453 0.0484
5 Discussion and Limitations
Our current evaluation focuses on low to medium dimensional PDE and ODE settings so that we can isolate the gains from the Chebyshev-based backbone, the internal feature re-injection and the rank-aware gating in a controlled regime. Extending AC-PKAN to genuinely high dimensional or chaotic systems such as Lorenz–96 would additionally require techniques for stable time integration, for controlling long-horizon error growth and for conditioning stiff gradients, which are outside the scope of this version Karimi and Paul (2010); Wang et al. (2022a); Maiocchi et al. (2024). Full layerwise Jacobian-rank profiling for ultra-deep stacks is infeasible on our 40 GB GPUs because backprop must retain large activations and Jacobian blocks. We list this as a scalability limitation and will perform ultra-deep evaluations once larger-memory hardware is available. We also rely on a standard AdamW optimiser and have not yet explored KAN-specific optimisation or structured pruning, which we view as promising directions for improving scalability and interpretability in future work.
@@ 6 Conclusion
We introduced AC-PKAN, a novel framework that enhances PINNs by integrating Cheby1KAN with traditional MLPs and augmenting them with internal and external attention mechanisms. This improves the model’s ability to capture complex patterns and dependencies, resulting in superior performance on challenging PDE tasks, including previous PINN failure modes and complex physical environments. The RGA mechanism enhances training stability and convergence by dynamically adjusting loss terms. Experimental results demonstrate that AC-PKAN consistently outperforms or matches state-of-the-art models like PINNsFormer, confirming its effectiveness in real-world engineering problems.
Appendix A Mathematical Proofs
A.1 Proof of Theorem 1
Lemma 1.
@@ A.1 Proof of Theorem 1
By the properties of subspace dimensions, it follows from equation 24 that
@@
Therefore, from equation 26 and equation 27, it follows that
@@
Combining equation 25 and equation 28, we obtain
@@
Applying the previously established result equation 29, we obtain
@@
By the inductive hypothesis equation 31, we have
@@
Therefore, substituting equation 34 into equation 33, we obtain
@@
By induction, the inequality equation 30 holds for all $n\geq 2$. This completes the proof. ∎
Theorem 4 (Single Cheb1KAN Layer Rank Constraint).
@@
To conclude that $N+1$ Chebyshev polynomials $T_{0}(x),T_{1}(x),\cdots,T_{N}(x)$ are linearly independent on interval [1,1][-1,1], we assume the opposite: there exist constants $c_{0},c_{1},\cdots,c_{N}$ such that:
$\forall x\in[-1,1],\qquad c_{0}T_{0}(x)+c_{1}T_{1}(x)+\cdots+c_{N}T_{N}(x)=0$ (39)
Since each $T_{k}(x)$ are a set of k-degree orthogonal polynomials according to 9, the left side is a polynomial of degree at most $N$. A non-zero polynomial of degree $N$ can have at most $N$ roots. However, the equation holds for all $x$ in [1,1][-1,1], which is an infinite set of points. Therefore, the polynomial must be the zero polynomial, implying $c_{0}=c_{1}=\cdots=c_{N}=0$.
@@
The key to this upper bound is that the basis function expansions for different input vectors are independent. Based on Equation 37, although the output dimension $d_{l+1}$ may be much smaller than $d_{l}\cdot(N+1)$, the final column space dimension is constrained by the following two factors:
@@
A.2 Proof of Theorem 2
Theorem 5 (Nonlinear Normalization Effect).